# Composing Concepts from Images and Videos via Concept-prompt Binding

### v1

@@ 1 Introduction
Visual concept composition aims to integrate different elements from images and videos into a single, coherent visual output. This process is a reflection of human artists’ creation: combining ingredients from various inspirations to form a brand new masterpiece [ipcomposer]. Consequently, it plays a fundamental role in visual creativity and filmmaking [film_survey]. With the rapid advancement of diffusion-based visual content generation models [sd3, labs2025flux1kontextflowmatching, easyanimate, cogvideox, opensora2, hunyuanvideo, hacohen2024ltxvideorealtimevideolatent, wan, data_synth_3dmm, musetalk, kong2025profashion, ifv2v], an increasing number of works [abdal2025zeroshotdynamicconceptpersonalization, dynamicconcept_singlevideo, tokenverse, videomage, dreamvideo, wang2025dualrealadaptivejointtraining, kwontweediemix, chen2025xverseconsistentmultisubjectcontrol, dalva2025lorashoptrainingfreemulticonceptimage, omg, conceptweaver] have been exploring the field of visual concept composition by exploiting the generative models’ strong capability of concept grounding and customization.
@@
Despite considerable efforts devoted to this field, challenges still remain in accurately extracting complex concepts from visual inputs and flexibly combining concepts from both images and videos. First, the capability to precisely extract specific concepts from various sources is of great significance for visual content creators. Nevertheless, existing mainstream methods [abdal2025zeroshotdynamicconceptpersonalization, dynamicconcept_singlevideo, videomage, dreamvideo, wang2025dualrealadaptivejointtraining, dalva2025lorashoptrainingfreemulticonceptimage, omg, conceptweaver] use either adapters like LoRA [lora] or learnable embeddings with explicit or implicit masks to realize concept selection, which fall short in decoupling complex concepts with occlusions and temporal alterations, and extracting non-object concepts such as styles. Second, it is a common practice to integrate different visual elements from both images and videos in the visual content creation process [film_survey]. However, previous works are confined to animating designated subjects from images with motion from videos [videomage, dreamvideo, wang2025dualrealadaptivejointtraining], without further exploration of flexibly combining various attributes (e.g., visual styles and lighting variations) from both images and videos. Although there has been recent effort on flexible concept composition [tokenverse] in the image domain, achieving universal visual concept composition for both images and videos remains an underexplored problem.
@@
To this end, we introduce Bind & Compose (BiCo), a one-shot method that enables flexible visual concept composition by binding visual concepts with the corresponding textual tokens, with satisfactory compatibility between image and video concepts (Fig. 1). Our method first leverages the powerful concept grounding capability [diffpng] of text-to-video (T2V) diffusion models [wan] to bind textual tokens with their corresponding visual concepts through one-shot training, achieving implicit decomposition without mask input. Then, concept composition is done through selecting any desired bound tokens from various sources and composing them into the final prompt tokens, which serves as the model condition. This paper mainly encompasses the following three technical contributions: First, to achieve reliable decomposition of complex visual concepts for flexible manipulation, we propose a hierarchical binder structure for the cross-attention mechanism [attention2017] in Diffusion Transformer (DiT) [dit] blocks to effectively encode visual concepts into corresponding textual tokens. When composing concepts from multiple sources, concept tokens in the target prompt are passed through different binders correspondingly to integrate visual features, enabling text-conditioned concept composition without explicit mask input. Second, to improve the accuracy of concept-token binding for more precise concept decomposition, we design a Diversify-and-Absorb Mechanism (DAM) that diversifies the one-shot prompts while retaining key concepts, and introduces an extra absorbent token during training to eliminate the impact of concept-irrelevant details. Third, to enhance the compatibility between image and video concepts during composition, we present a Temporal Disentanglement Strategy (TDS) that decouples the training process of video concepts into two stages. In the first stage, the binders are trained with individual frames without temporal concepts, which aligns with the training setting of image concepts. In the second stage that trains the binders on videos, we adopt a dual-branch binder structure to better cater to temporal concepts while inheriting knowledge from the first stage.
@@ 2 Related Work
T2V Diffusion Models. The emergence of diffusion models [ddpm, ddim, ldm] has significantly advanced the realm of visual content generation. Recently, DiT [dit] has become the de facto standard of the denoising model’s architecture, surpassing U-Net [unet] with its strong scaling capability [kaplan2020scalinglawsneurallanguage] and flexibility for integrating multi-modal conditions [sd3]. Flow Matching [flowmatching, rectifiedflow] introduces a new linear paradigm to transit between the Gaussian distribution and the target distribution, improving the theoretical properties and simplifying the conceptual framework. Based on these works, a number of T2V diffusion models [easyanimate, cogvideox, opensora2, hunyuanvideo, hacohen2024ltxvideorealtimevideolatent, wan] have emerged with text-to-video cross-attention or joint attention for conditioning. Despite these methods achieving satisfactory quality and consistency of generation, they are designed for general T2V generation, with limited support for personalization and concept composition.
@@
Video Personalization. Video personalization aims to integrate the appearance or motion of a designated object into a pre-trained video generation model, enabling the model to reproduce these properties when generating with other prompts. Building upon the progress in the image domain [dreambooth, iclora, diffusion_self_distillation], several approaches handle the temporal consistency problem by adding LoRAs [lora] to the temporal layers of T2V models [customttt, he2024idanimatorzeroshotidentitypreservinghuman] or learning the motion embeddings from reference videos [newmove, dreamvideo, motiondirector]. Set-and-Sequence [set_and_sequence] enables the simultaneous learning of both appearance and motion from a single video by designing the spatio-temporal weight space within the LoRA architecture. Grid-LoRA [gridlora] further enables reusable video personalization by introducing a grid-based LoRA system that spatially organizes input and output. However, we cannot accurately designate the concept to extract and the way the concepts are combined. The number and type of inputs are also confined, limiting the flexibility of concept composition.
Visual Concept Composition. Composing multiple visual elements from images and videos into a coherent output remains a challenging task. There have been early explorations in decomposing image concepts [breakascene, conceptexpress, inspirationtree]. Break-A-Scene [breakascene] relies on explicit mask inputs to achieve concept decomposition, which limits its availability to common users and its ability to extract non-object concepts. Other methods [conceptexpress, inspirationtree] extract multiple concepts from a single image by jointly learning several tokens, each corresponding to a visual concept. However, the content learned by each token is unpredictable. To achieve concept composition in the image domain, existing works either use explicit spatial conditioning [omg, yang2024loracomposerleveraginglowrankadaptation, conceptweaver], which falls short in overlapping or non-object concepts, or fuse multiple LoRAs [mixofshow, modular_customization, ziplora], which restricts the type and number of concepts to compose or requires joint optimization among all source images. TokenVerse [tokenverse] learns a modulation term for each text token to achieve prompt-controlled concept composition. Despite enhanced flexibility, it relies on text-conditioned modulation architectures in DiT [dit] models, limiting its universality to modern T2V models [wan, meituanlongcatteam2025longcatvideotechnicalreport]. To extend concept composition to handle videos, previous methods [dreamvideo, wang2025dualrealadaptivejointtraining, videomage] incorporate dedicated designs to decouple appearance and motion, supporting only the composition of subjects from images and motions from videos. BiCo simultaneously enables complex concept decomposition (non-object concepts and multiple concepts from a single input) and flexible concept composition (selective composition via prompts and composing image and video concepts), offering endless possibilities for visual creators.
@@ 3.1 Overview
Given $M$ concept images or videos $\{\bm{V}^{j}_{c}\}_{j=1}^{M}$ with their corresponding textual prompt tokens $\{\bm{p}^{j}_{c}\}_{j=1}^{M}$, BiCo aims at composing the visual concepts from the inputs according to the designated prompt $\bm{p}_{d}$ to generate a coherent visual output $\bm{V}$. As illustrated in Fig. 2, it first learns each text token’s corresponding visual appearance or motion via a light-weight binder module for each visual input, and then combines tokens from different source images or videos to generate a target video that composes the individual concepts. Specifically, during concept binding, a binder structure attached to a DiT-based T2V model [wan] is utilized to encode the correspondence between visual concepts and textual tokens through one-shot training on different inputs $\{\bm{V}^{j}_{c}\}_{j=1}^{M}$ and $\{\bm{p}^{j}_{c}\}_{j=1}^{M}$ respectively. When integrating concepts from various sources, different parts of the designated prompt $\bm{p}_{d}$ representing visual concepts are passed through their corresponding binders to compose the updated prompt $\bm{p}_{u}$, which contains visual concept information and is then fed into DiT blocks to serve as the condition for the composed visual output.
Preliminary: Text Conditioning in T2V Models. Current mainstream T2V models [easyanimate, cogvideox, opensora2, hunyuanvideo, hacohen2024ltxvideorealtimevideolatent, wan] adopt the DiT [dit] architecture with tens of blocks to predict the denoising vector. Each DiT block contains attention layers, an MLP, and a modulation mechanism for the timestep condition. To achieve text conditioning, a prevalent method is to insert a cross-attention layer in each DiT block, which takes the latent tokens $\bm{x}_{in}$ as queries and the textual prompt tokens $\bm{p}$ as keys and values:
@@ 3.2 Hierarchical Binder Structure
To fully exploit the powerful capability of visual-text association within T2V models for accurate decomposition of complex visual concepts, we attach binders to DiT cross-attention conditioning layers to encode visual concepts into corresponding prompt tokens. Since DiT blocks have distinct behaviors during the denoising process [cache_me_if_you_can], a hierarchical binder structure is designed with a global binder for the overall association and per-block binders for tailored association (Fig. 3). Specifically, each binder $f(\cdot)$ consists of an MLP with a zero-initialized learnable scaling factor $\gamma$ in a residual style, and takes the prompt tokens $\bm{p}$ as input:
@@
Two-stage Inverted Training Strategy. Recent studies point out that the denoising process of diffusion models is divided into several stages with different functions [autodiffusion, perception_prioritized_training]. It has been discovered that prioritizing the training on higher noise levels yields better performance [perception_prioritized_training]. To this end, we utilize a two-stage inverted training strategy to enhance the optimization process for hierarchical binders. Specifically, we define a noise level threshold $\alpha$ to separate high and low noise levels. In the first stage, we only train the global binder with the probability of $\alpha$ for high noise levels ($\geq\alpha$) and the probability of $1-\alpha$ for low noise levels ($\alpha$). This setting emphasizes the high noise levels and reduces the optimization steps on low noise levels. In the second training stage, both global and per-block binders are trained without inverting the probability of noise levels.
@@ 3.3 Diversify-and-Absorb Mechanism (DAM)
Establishing accurate concept-token bindings is a notable challenge, especially in one-shot cases. To enable precise association between concepts and prompt tokens in binders, we introduce DAM, which takes advantage of the powerful visual comprehension and reasoning capability of Vision-language Models (VLMs) [bai2025qwen25vltechnicalreport] to diversify concept prompts while retaining the key conceptual words unchanged during the concept binding process. As shown in Fig. 4, the prompt diversification process is divided into two stages: key concept extraction and spatiotemporal concept composition. In the key concept extraction stage, the VLM is asked to extract critical concepts from the input image or video and divide them into spatial and temporal concepts. During spatiotemporal concept composition, the VLM composes the extracted concepts into a designated number of diverse prompts with the visual input reference. For images and the first-stage training of videos with a focus on spatial concepts (detailed in §3.4), only spatial concepts are used to form the full prompt. For the second-stage training of videos, the VLM uses both spatial and temporal concepts to generate the complete prompt.
@@ 3.4 Temporal Disentanglement Strategy (TDS)
The ability to compose concepts from both images and videos is of great significance to visual content creators. However, significant temporal heterogeneity exists between images and videos [spatial_temporal_causal_inference], especially the temporal domain shift caused by the absence of motion in images. This hinders compatibility when directly composing concepts from both sources. To enable flexible composition of image and video concepts with satisfactory quality, we devise TDS, which aligns the learning paradigm of images and videos by decoupling the training process for video concepts into two stages. In the first stage, we train the binders on individual video frames without temporal concepts in the input prompt. This setting remains the same as the training setting of image concepts, with a focus on binding spatial concepts. In the second stage that trains the binders on full videos and complete prompts with temporal concepts, we adopt a dual-branch binder structure to decouple the learning of spatial and temporal concepts and inherit the knowledge from the first stage. Specifically, we extend the MLP in the original binder with an extra temporal MLP branch $\texttt{MLP}_{t}$, and then fuse them with a learnable gating module $g(\cdot)$:
@@ 4.1 Implementation Details
We select Wan2.1-T2V-1.3B [wan] as the base model to apply BiCo. The MLP structure in binders consists of two linear layers with layer normalization [ba2016layernormalization] and GELU [gelu] activation. The binders are trained with a learning rate of $1.0\times 10^{-4}$ for $2400$ iterations per stage. The noise level threshold $\alpha$ in §3.2 is set to $0.875$. We set the length for composed videos to $81$ frames during inference. All other hyperparameters remain the same as Wan2.1 [wan]. Experiments are conducted on NVIDIA RTX 4090 GPUs.
@@ 4.2 Comparisons to Prior Works
To demonstrate the superiority of BiCo over existing visual concept composition works, we conduct quantitative and qualitative comparisons with 4 representative methods: Textual Inversion (Text-Inv) [textualinversion], DreamBooth-LoRA (DB-LoRA) [dreambooth], DreamVideo [dreamvideo], and DualReal [wang2025dualrealadaptivejointtraining]. We adapt Text-Inv and DB-LoRA on the same T2V model [wan] as BiCo to support video concepts. Since existing methods that support both images and videos only take one image (subject) and one video (motion) as input, we limit our comparisons to composing concepts from one image and one video for fair comparisons in this section.

### v2

@@ 1 Introduction
Visual concept composition aims to integrate different elements from images and videos into a single, coherent visual output. This process is a reflection of human artists’ creation: combining ingredients from various inspirations to form a brand new masterpiece [15]. Consequently, it plays a fundamental role in visual creativity and filmmaking [62]. With the rapid advancement of diffusion-based visual content generation models [16, 35, 58, 61, 42, 29, 20, 54, 40, 63, 31, 30], an increasing number of works [1, 3, 18, 26, 56, 55, 34, 11, 14, 32, 33] have been exploring the field of visual concept composition by exploiting the generative models’ strong capability of concept grounding and customization.
@@
Despite considerable efforts devoted to this field, challenges still remain in accurately extracting complex concepts from visual inputs and flexibly combining concepts from both images and videos. First, the capability to precisely extract specific concepts from various sources is of great significance for visual content creators. Nevertheless, existing mainstream methods [1, 3, 26, 56, 55, 14, 32, 33] use either adapters like LoRA [25] or learnable embeddings with explicit or implicit masks to realize concept selection, which fall short in decoupling complex concepts with occlusions and temporal alterations, and extracting non-object concepts such as styles. Second, it is a common practice to integrate different visual elements from both images and videos in the visual content creation process [62]. However, previous works are confined to animating designated subjects from images with motion from videos [26, 56, 55], without further exploration of flexibly combining various attributes (e.g., visual styles and lighting variations) from both images and videos. Although there has been recent effort on flexible concept composition [18] in the image domain, achieving universal visual concept composition for both images and videos remains an underexplored problem.
@@
To this end, we introduce Bind & Compose (BiCo), a one-shot method that enables flexible visual concept composition by binding visual concepts with the corresponding textual tokens, with satisfactory compatibility between image and video concepts (Fig. 1). Our method first leverages the powerful concept grounding capability [59] of text-to-video (T2V) diffusion models [54] to bind textual tokens with their corresponding visual concepts through one-shot training, achieving implicit decomposition without mask input. Then, concept composition is done through selecting any desired bound tokens from various sources and composing them into the final prompt tokens, which serves as the model condition. This paper mainly encompasses the following three technical contributions: First, to achieve reliable decomposition of complex visual concepts for flexible manipulation, we propose a hierarchical binder structure for the cross-attention mechanism [52] in Diffusion Transformer (DiT) [41] blocks to effectively encode visual concepts into corresponding textual tokens. When composing concepts from multiple sources, concept tokens in the target prompt are passed through different binders correspondingly to integrate visual features, enabling text-conditioned concept composition without explicit mask input. Second, to improve the accuracy of concept-token binding for more precise concept decomposition, we design a Diversify-and-Absorb Mechanism (DAM) that diversifies the one-shot prompts while retaining key concepts, and introduces an extra absorbent token during training to eliminate the impact of concept-irrelevant details. Third, to enhance the compatibility between image and video concepts during composition, we present a Temporal Disentanglement Strategy (TDS) that decouples the training process of video concepts into two stages. In the first stage, the binders are trained with individual frames without temporal concepts, which aligns with the training setting of image concepts. In the second stage that trains the binders on videos, we adopt a dual-branch binder structure to better cater to temporal concepts while inheriting knowledge from the first stage.
@@ 2 Related Work
T2V Diffusion Models. The emergence of diffusion models [24, 50, 46] has significantly advanced the realm of visual content generation. Recently, DiT [41] has become the de facto standard of the denoising model’s architecture, surpassing U-Net [47] with its strong scaling capability [28] and flexibility for integrating multi-modal conditions [16]. Flow Matching [37, 38] introduces a new linear paradigm to transit between the Gaussian distribution and the target distribution, improving the theoretical properties and simplifying the conceptual framework. Based on these works, a number of T2V diffusion models [58, 61, 42, 29, 20, 54] have emerged with text-to-video cross-attention or joint attention for conditioning. Despite these methods achieving satisfactory quality and consistency of generation, they are designed for general T2V generation, with limited support for personalization and concept composition.
@@
Video Personalization. Video personalization aims to integrate the appearance or motion of a designated object into a pre-trained video generation model, enabling the model to reproduce these properties when generating with other prompts. Building upon the progress in the image domain [48, 27, 9], several approaches handle the temporal consistency problem by adding LoRAs [25] to the temporal layers of T2V models [8, 22] or learning the motion embeddings from reference videos [39, 56, 64]. Set-and-Sequence [4] enables the simultaneous learning of both appearance and motion from a single video by designing the spatio-temporal weight space within the LoRA architecture. Grid-LoRA [2] further enables reusable video personalization by introducing a grid-based LoRA system that spatially organizes input and output. However, we cannot accurately designate the concept to extract and the way the concepts are combined. The number and type of inputs are also confined, limiting the flexibility of concept composition.
Visual Concept Composition. Composing multiple visual elements from images and videos into a coherent output remains a challenging task. There have been early explorations in decomposing image concepts [5, 21, 53]. Break-A-Scene [5] relies on explicit mask inputs to achieve concept decomposition, which limits its availability to common users and its ability to extract non-object concepts. Other methods [21, 53] extract multiple concepts from a single image by jointly learning several tokens, each corresponding to a visual concept. However, the content learned by each token is unpredictable. To achieve concept composition in the image domain, existing works either use explicit spatial conditioning [32, 60, 33], which falls short in overlapping or non-object concepts, or fuse multiple LoRAs [19, 44, 49], which restricts the type and number of concepts to compose or requires joint optimization among all source images. TokenVerse [18] learns a modulation term for each text token to achieve prompt-controlled concept composition. Despite enhanced flexibility, it relies on text-conditioned modulation architectures in DiT [41] models, limiting its universality to modern T2V models [54, 51]. To extend concept composition to handle videos, previous methods [56, 55, 26] incorporate dedicated designs to decouple appearance and motion, supporting only the composition of subjects from images and motions from videos. BiCo simultaneously enables complex concept decomposition (non-object concepts and multiple concepts from a single input) and flexible concept composition (selective composition via prompts and composing image and video concepts), offering endless possibilities for visual creators.
@@ 3.1 Overview
Given $M$ concept images or videos $\{\bm{V}^{j}_{c}\}_{j=1}^{M}$ with their corresponding textual prompt tokens $\{\bm{p}^{j}_{c}\}_{j=1}^{M}$, BiCo aims at composing the visual concepts from the inputs according to the designated prompt $\bm{p}_{d}$ to generate a coherent visual output $\bm{V}$. As illustrated in Fig. 2, it first learns each text token’s corresponding visual appearance or motion via a light-weight binder module for each visual input, and then combines tokens from different source images or videos to generate a target video that composes the individual concepts. Specifically, during concept binding, a binder structure attached to a DiT-based T2V model [54] is utilized to encode the correspondence between visual concepts and textual tokens through one-shot training on different inputs $\{\bm{V}^{j}_{c}\}_{j=1}^{M}$ and $\{\bm{p}^{j}_{c}\}_{j=1}^{M}$ respectively. When integrating concepts from various sources, different parts of the designated prompt $\bm{p}_{d}$ representing visual concepts are passed through their corresponding binders to compose the updated prompt $\bm{p}_{u}$, which contains visual concept information and is then fed into DiT blocks to serve as the condition for the composed visual output.
Preliminary: Text Conditioning in T2V Models. Current mainstream T2V models [58, 61, 42, 29, 20, 54] adopt the DiT [41] architecture with tens of blocks to predict the denoising vector. Each DiT block contains attention layers, an MLP, and a modulation mechanism for the timestep condition. To achieve text conditioning, a prevalent method is to insert a cross-attention layer in each DiT block, which takes the latent tokens $\bm{x}_{in}$ as queries and the textual prompt tokens $\bm{p}$ as keys and values:
@@ 3.2 Hierarchical Binder Structure
To fully exploit the powerful capability of visual-text association within T2V models for accurate decomposition of complex visual concepts, we attach binders to DiT cross-attention conditioning layers to encode visual concepts into corresponding prompt tokens. Since DiT blocks have distinct behaviors during the denoising process [57], a hierarchical binder structure is designed with a global binder for the overall association and per-block binders for tailored association (Fig. 3). Specifically, each binder $f(\cdot)$ consists of an MLP with a zero-initialized learnable scaling factor $\gamma$ in a residual style, and takes the prompt tokens $\bm{p}$ as input:
@@
Two-stage Inverted Training Strategy. Recent studies point out that the denoising process of diffusion models is divided into several stages with different functions [36, 13]. It has been discovered that prioritizing the training on higher noise levels yields better performance [13]. To this end, we utilize a two-stage inverted training strategy to enhance the optimization process for hierarchical binders. Specifically, we define a noise level threshold $\alpha$ to separate high and low noise levels. In the first stage, we only train the global binder with the probability of $\alpha$ for high noise levels ($\geq\alpha$) and the probability of $1-\alpha$ for low noise levels ($\alpha$). This setting emphasizes the high noise levels and reduces the optimization steps on low noise levels. In the second training stage, both global and per-block binders are trained without inverting the probability of noise levels.
@@ 3.3 Diversify-and-Absorb Mechanism (DAM)
Establishing accurate concept-token bindings is a notable challenge, especially in one-shot cases. To enable precise association between concepts and prompt tokens in binders, we introduce DAM, which takes advantage of the powerful visual comprehension and reasoning capability of Vision-language Models (VLMs) [7] to diversify concept prompts while retaining the key conceptual words unchanged during the concept binding process. As shown in Fig. 4, the prompt diversification process is divided into two stages: key concept extraction and spatiotemporal concept composition. In the key concept extraction stage, the VLM is asked to extract critical concepts from the input image or video and divide them into spatial and temporal concepts. During spatiotemporal concept composition, the VLM composes the extracted concepts into a designated number of diverse prompts with the visual input reference. For images and the first-stage training of videos with a focus on spatial concepts (detailed in §3.4), only spatial concepts are used to form the full prompt. For the second-stage training of videos, the VLM uses both spatial and temporal concepts to generate the complete prompt.
@@ 3.4 Temporal Disentanglement Strategy (TDS)
The ability to compose concepts from both images and videos is of great significance to visual content creators. However, significant temporal heterogeneity exists between images and videos [12], especially the temporal domain shift caused by the absence of motion in images. This hinders compatibility when directly composing concepts from both sources. To enable flexible composition of image and video concepts with satisfactory quality, we devise TDS, which aligns the learning paradigm of images and videos by decoupling the training process for video concepts into two stages. In the first stage, we train the binders on individual video frames without temporal concepts in the input prompt. This setting remains the same as the training setting of image concepts, with a focus on binding spatial concepts. In the second stage that trains the binders on full videos and complete prompts with temporal concepts, we adopt a dual-branch binder structure to decouple the learning of spatial and temporal concepts and inherit the knowledge from the first stage. Specifically, we extend the MLP in the original binder with an extra temporal MLP branch $\texttt{MLP}_{t}$, and then fuse them with a learnable gating module $g(\cdot)$:
@@ 4.1 Implementation Details
We select Wan2.1-T2V-1.3B [54] as the base model to apply BiCo. The MLP structure in binders consists of two linear layers with layer normalization [6] and GELU [23] activation. The binders are trained with a learning rate of $1.0\times 10^{-4}$ for $2400$ iterations per stage. The noise level threshold $\alpha$ in §3.2 is set to $0.875$. We set the length for composed videos to $81$ frames during inference. All other hyperparameters remain the same as Wan2.1 [54]. Experiments are conducted on NVIDIA RTX 4090 GPUs.
@@ 4.2 Comparisons to Prior Works
To demonstrate the superiority of BiCo over existing visual concept composition works, we conduct quantitative and qualitative comparisons with 4 representative methods: Textual Inversion (Text-Inv) [17], DreamBooth-LoRA (DB-LoRA) [48], DreamVideo [56], and DualReal [55]. We adapt Text-Inv and DB-LoRA on the same T2V model [54] as BiCo to support video concepts. Since existing methods that support both images and videos only take one image (subject) and one video (motion) as input, we limit our comparisons to composing concepts from one image and one video for fair comparisons in this section.